OpenStudy (anonymous):

Help... #3 the answer is correct, but leave it as a square root. #4 you have one set of numbers listed, but you did not identify them as a domain or range. There should be two sets of numbers listed. #10 double check the exponent for y #11 & 15 where are these questions? ALL level 2 and 3 questions have to be answered before I can give you a grade. #13 What kind of function is this? #14 shifted to the left how many units?

OpenStudy (anonymous):

I have all the anwsers, I just need the stuff mentioned above... thanks

OpenStudy (anonymous):

10sqr3+2sqr3=12sqr3

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

TheMind is saying this \[\Large 10\sqrt{3}+2\sqrt{3} = 12\sqrt{3}\]

OpenStudy (anonymous):

For what number?

jimthompson5910 (jim_thompson5910):

#3

OpenStudy (anonymous):

what about the others?

jimthompson5910 (jim_thompson5910):

# 4 Domain: {2, 8, 0, –3} Range: {5,17,0,–5}

jimthompson5910 (jim_thompson5910):

What did you get for #10?

OpenStudy (anonymous):

correct,, jim_thompson5910

OpenStudy (anonymous):

umm you said 10?

jimthompson5910 (jim_thompson5910):

yes, what did you get for #10?

jimthompson5910 (jim_thompson5910):

no i get a server error

jimthompson5910 (jim_thompson5910):

i see what went wrong, y^2 divided by y^-5 is y^(2-(-5)) = y^(2+5) = y^7 So (y^7)^3 = y^(7*3) = y^(21)

OpenStudy (anonymous):

10?

jimthompson5910 (jim_thompson5910):

for #10 yes

OpenStudy (anonymous):

what aBOUT 11,13, AND 15

jimthompson5910 (jim_thompson5910):

# 11 what did you get? yes or no ------------------------------- # 12 It's not a function since it fails the vertical line test (a vertical line can be drawn where it passes through more than one point) ------------------------------- # 13 Correct

jimthompson5910 (jim_thompson5910):

# 15 what did you get for this one?

OpenStudy (anonymous):

its on the paper i think

jimthompson5910 (jim_thompson5910):

I don't see #15

OpenStudy (anonymous):

what function is 13, and what about 14

jimthompson5910 (jim_thompson5910):

# 13 the parent function is \(\Large f(x) = x^2\)

jimthompson5910 (jim_thompson5910):

for #14, the parent function is \(\Large f(x) = \sqrt{x}\)

OpenStudy (anonymous):

#14 shifted to the left how many units?

jimthompson5910 (jim_thompson5910):

2 because the endpoint of \(\Large f(x) = \sqrt{x}\) is at (0,0)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

@jim_thompson5910 #14. What you've told me about the parent function is true, but now you need to tell me how the function in the question has been moved, compared to the parent function

jimthompson5910 (jim_thompson5910):

you already said it: shift the parent function 2 units to the left

OpenStudy (anonymous):

@jim_thompson5910 Only 2 units? Look a little closer...

jimthompson5910 (jim_thompson5910):

oh you caught it before I did, it's 3, not 2

OpenStudy (anonymous):

haha yeah.... I was looking at tit, i was like "wait"